# Optimal Transportation by Orthogonal Coupling Dynamics

### v1

@@ 3.2. Elliptic Distributions
which results from $p_{t}$ being Gaussian, see [23]. The matrix-Riccati equation (35) admits many stationary solutions, since any bijection between $\mu$ and $\nu$ is a stationary solution of (2.1), as discussed in [23]. However, as we justified in Proposition 2.2, starting with independent samples, i.e. $\mathbb{E}[X_{0}Y_{0}]=0$, $J_{t}$ remains symmetric positive semi-definite and its trace increases monotonically. This restricts the stationary solution to the one associated with the optimal transport. More precisely, we have
$\dot{X}_{t}=Y_{t}-J_{t}^{T}\Sigma_{\mu}^{-1}X_{t}\ \ \ \textrm{and}\ \ \ \dot{Y}_{t}=X_{t}-J_{t}\Sigma_{\nu}^{-1}Y_{t}$ (37)
with the fixed point
$Y_{s}=J_{s}^{T}\Sigma_{\mu}^{-1}X_{s}\ \ \ \textrm{and}\ \ \ X_{s}=J_{s}\Sigma_{\nu}^{-1}Y_{s}\ .$ (38)
Given that $J_{s}$ is symmetric positive semi-definite (Proposition 2.2), each of $X_{s}$ or $Y_{s}$ can be expressed as gradient of a convex function of the other (similar to Eq. (34)). For the one-dimensional case, Eq. (35) takes a simpler form. Let $n=1$ and suppose the variances are $\sigma^{2}_{\mu}$ and $\sigma^{2}_{\nu}$. Define the correlation coefficient $\kappa_{t}\mathrel{\mathop{:}}=J_{t}/(\sigma_{\mu}\sigma_{\nu})$. Therefore, the Riccati equation (35) reduces to
$\partial_{t}\kappa_{t}$ $=$ $\left(\frac{\sigma_{\mu}^{2}+\sigma_{\nu}^{2}}{\sigma_{\mu}\sigma_{\nu}}\right)\left(1-\kappa_{t}^{2}\right)\ .$ (39)
Starting from independent samples, this nonlinear ODE gives a monotone increase of $\kappa_{t}$ from zero to the stationary solution $\kappa_{t}\to 1$ as $t\to\infty$, which is consistent with the solution of the optimal transport problem (note that here the Monge map is given by $\hat{T}(X)=\sigma_{\nu}/\sigma_{\mu}X$ leading to unity correlation coefficient).
@@ 4. Numerical Algorithm through Opinion Dynamics
Opinion dynamics at its core tries to answer how different agents shape their opinions in contact with each-other [37]. While seemingly unrelated to the Monge-Kantorovich problem, in this section, we leverage analogy between opinion dynamics and OCD to provide a novel, yet simple, numerical algorithm for the latter. The main idea is to first, split the data points into a certain number of clusters. Next, the conditional expectation is estimated in each cluster using an appropriate hypothesis class.
@@ 4.1. Analogy with Opinion Dynamics
Consider a population of agents indexed by $i\in\{1,...,N\}$. Suppose each person $i$ has certain two-dimensional position $P_{i}(t)\in\mathbb{R}^{2}$ as their opinion in time $t$. In the standard form of consensus dynamics, e.g. see [44], these positions relax towards a global consensus via
@@
    The positions are perfectly tied to each-other and $\alpha$ becomes the exchange matrix $J$.
@@
$\dot{P}_{X,i}=\frac{1}{N_{\mathrm{bar}{I}^{X_{i},\epsilon}}}\sum_{j\in\mathrm{bar}{I}^{X_{i},\epsilon}}(P_{Y,i}(t)-P_{Y,j}(t))\ \ \textrm{and}\ \ \dot{P}_{Y,i}=\frac{1}{N_{\mathrm{bar}{I}^{Y_{i},\epsilon}}}\sum_{j\in\mathrm{bar}{I}^{Y_{i},\epsilon}}(P_{X,i}(t)-P_{X,j}(t))\ ,$ (42)
@@
which is OCD (2.1) for $L^{2}$-cost. From technical point of view, the discrete version is justified due to the approximation of conditional expectation by a discrete kernel [46]. Through the lens of opinion dynamics, the OCD describes a dynamics which is local in a sense that each agent is only influenced by agents in its vicinity, yet as discussed before, it is closely linked to the global optimization problem (1).
@@ 4.2. Non-parametric Monte-Carlo Algorithm
Starting with $N_{p}$ i.i.d. samples $\hat{X}_{i,t_{0}}\sim\mu$ and $\hat{Y}_{i,t_{0}}\sim\nu$, we update the state of each sample according to discretized form of OCD. It is important to note that here we are focusing on an approximate solution of the optimization problem (1) over atom-less distributions $\mu$ and $\nu$. An alternative interpretation, where the distributions are assumed to be sum of diracs (assignment problem) is not pursued here (see [62]). Therefore throughout what follows, we assume that there exists an underlying continuous map between the two measure spaces. Three hyper-parameters control the discretization error: number of samples $N_{p}$, time step size $\Delta t$, and cluster cut-off $\epsilon$. Once these values are fixed, relying on Euler’s scheme (see Runge-Kutta’s scheme in Sec.7.2 of Supplementary Information), each sample is updated according to
@@
    OCD-piecewise constant. The estimator is constant in each cluster (equivalent to using square kernel in the estimation of the conditional expectation). The expressions for $\hat{\mathcal{K}}^{n}_{X,Y}(.)$ are given in Eqs. (76)-(77) in Supplementary Information.
@@
    OCD-piecewise linear. The estimator follows linear regression (accounts for using Gaussian kernel in the estimation of the conditional expectation). In other words here we look for $L^{2}$ projection of $\nabla_{x,y}c$ onto the span of linear functions [36]. The final expressions are given in Eqs. (85)-(86) in Supplementary Information.
The overview of the algorithm is given in Algorithm 1 in Supplementary Information. The convergence of the algorithm, in general, is controlled by $N_{p},\Delta t,\epsilon$. Assuming that the OCD initialized by samples of $\mu,\nu$ is well-posed, it is reasonable to consider fixed-order convergence with respect to $\Delta t$ (e.g. first-order for Euler time integration, see [50]). Also the error due to finite $N_{p}$ is expected to behave similar to conventional Monte-Carlo schemes [38]. However, the error dependency on the cut-off value $\epsilon$ is more delicate. Two asymptotic limits of $\epsilon\to 0$ and $\epsilon\to\infty$ can be intuitively understood. The limit $\epsilon\to 0$ accounts for no interaction between agents, i.e. each point forms its own cluster. Therefore the points remain frozen in time. On the other hand, $\epsilon\to\infty$ gives rise to perfect interaction among all agents, forming one giant cluster. In this case, the dynamics converges to the average value among all sample points, for the 1st order approximation, and towards linear regression, for the 2nd order case. Obviously, the optimal value of $\epsilon$ needs to be tuned in-between, and is expected to depend on $N_{p}$. The transition between frozen dynamics in $\epsilon\to 0$ and homogenization $\epsilon\to\infty$ suggests a phase transition type behaviour, resulting from $\epsilon$. The transition point can serve as a proxy for the optimal cut-off value, as discussed in Sec. 7.3 in Supplementary Information.
Once the pairing between $\hat{X}_{i}$ and $\hat{Y}_{i}$ reaches its stationary state (according to a stopping criterion), the sampled points $(\hat{X}_{i}^{s},\hat{Y}_{i}^{s})_{i\in\{1,...,N_{p}\}}$ can be employed to estimate the Monge map $\hat{T}()$, using a suitable hypothesis class. We train a Neural Network (NN) with a few layers and $\tanh(.)$ as the activation function to approximate the map and denote it by $M_{X\rightarrow Y}$. As the loss function, we simply consider the expectation of $L^{2}$ point-wise error between NN’s output $M_{X\rightarrow Y}(\hat{X}^{*})$ and $\hat{Y}^{*}$, i.e. $\mathrm{Loss}=\sum_{i=1}^{N_{p}}\lVert M_{X\rightarrow Y}(\hat{X}_{i}^{*})-\hat{Y}_{i}^{*}\rVert_{2}^{2}/N_{p}$.
@@ 4.3. Computational Complexity
For a given time step (iteration), the main computational task of the proposed OCD algorithm is evaluation of the conditional expectation, whose estimation relies on forming clusters by finding neighbours of each point in a given distribution. For the latter, though non-exclusively, the Ball Tree method [47] scales efficiently i.e. $\mathcal{O}(N_{p}\log N_{p})$ with number of particles $N_{p}$. Furthermore, the cost of OCD scales linearly with $n=\dim(X)=\dim(Y)$, since the number of ODEs is $2n$ and number of operations in order to compute the distance between each two particles is $\mathcal{O}(n)$. In terms of memory consumption, since OCD, unlike linear programming, does not require computation of the distance matrix, its memory requirement is $\mathcal{O}(N_{p})$. In order to estimate the total computational cost of OCD, including number of iterations required to reach stationary state, we conducted numerical tests as shown in Fig. 2, where OCD runs between two normal distributions. The flops and storage requirements confirm our analysis on complexity and memory consumption. However, it is evident that OCD requires significantly more operations compared to the reference Earth Mover’s Distance (EMD) method [29], if relatively small number of sample points (e.g. 100) is considered. This is mainly due to the fact that while the number of operations scales efficiently, i.e. with $\mathcal{O}(N_{p}\log N_{p})$, the pre-factor which depends on the number of iterations required to reach stationary condition becomes dominant, if $N_{p}$ is relatively small.
@@
(a) $\dim(X)=3$ (b) $N_{p}=800$
Figure 2. Execution time (top) and memory consumption (down) of OCD-piecewise linear algorithm with Runge Kuta 4th order solver with $\Delta t=0.1$ against Earth Mover’s Distance (EMD) method [29] in learning the map between samples of $\mu=N(0,I)$ and $\nu=N(1,I)$ in (a) $\dim(X)=3$ for a range of $N_{p}$ and (b) $\dim(X)=1,...,8$ with $N_{p}=800$. Here we scale $\epsilon$ with a rule of thumb $\varepsilon=3d{N_{p}}^{-1/4}/4$.

### v2

@@ 4. Characterization in Gaussian Setting
which results from $p_{t}$ being Gaussian, see [conforti2023projected]. The matrix-Riccati equation (44) admits many stationary solutions, since any bijection between $\mu$ and $\nu$ is a stationary solution of (14), as discussed in [conforti2023projected]. However, as we justified in Proposition 2.2, starting with independent samples, i.e. $\mathbb{E}[X_{0}Y_{0}]=0$, $J_{t}$ remains symmetric positive semi-definite and its trace increases monotonically. This restricts the stationary state and results in the OT solution if $\Sigma_{\mu}$ and $\Sigma_{\nu}$ commute.
Proposition 4.1 (OCD recovers OT for commuting Gaussian covariances).
Let
$\mu=\mathcal{N}(0,\Sigma_{\mu}),\qquad\nu=\mathcal{N}(0,\Sigma_{\nu}),$
with $\Sigma_{\mu},\Sigma_{\nu}\succ 0$, and let $(X_{t},Y_{t})$ evolve according to the $L^{2}$-OCD dynamics. If the covariances commute, i.e.
$\Sigma_{\mu},\Sigma_{\nu}]=0,$
then $L^{2}$-OCD converges to the optimal transport correlation, namely
$J_{t}\;\longrightarrow\;J_{\infty}=\Sigma_{\mu}^{1/2}\big(\Sigma_{\mu}^{1/2}\Sigma_{\nu}\Sigma_{\mu}^{1/2}\big)^{1/2}\Sigma_{\mu}^{1/2}=J_{\mathrm{OT}}.$}
Proof.
When $\Sigma_{\mu},\Sigma_{\nu}]=0$, the two matrices are simultaneously diagonalizable:
$\Sigma_{\mu}=Q\,\mathrm{diag}(\sigma_{\mu,i})\,Q^{\top},\qquad\Sigma_{\nu}=Q\,\mathrm{diag}(\sigma_{\nu,i})\,Q^{\top},$
for some orthogonal $Q$. Under this change of coordinates, (44) decouples component-wise
$\dot{j}_{i}=\sigma_{\mu,i}+\sigma_{\nu,i}-(\sigma_{\mu,i}^{-1}+\sigma_{\nu,i}^{-1})\,j_{i}^{2},\qquad j_{i}(0)=0.$
By direct computation we find that the equilibrium (given $j_{i}(0)=0$) is
$j_{i}^{\star}=\sqrt{\sigma_{\mu,i}\sigma_{\nu,i}}\ .$
Hence $J_{t}\to J_{\infty}=Q\,\mathrm{diag}(\sqrt{\sigma_{\mu,i}\sigma_{\nu,i}})\,Q^{\top}.$ On the other hand, the Gaussian OT cross–correlation is
$J_{\mathrm{OT}}=\Sigma_{\mu}^{1/2}(\Sigma_{\mu}^{1/2}\Sigma_{\nu}\Sigma_{\mu}^{1/2})^{1/2}\Sigma_{\mu}^{1/2},$
see e.g. [gelbrich1990formula]. When $\Sigma_{\mu}$ and $\Sigma_{\nu}$ commute, we get
$J_{\mathrm{OT}}=Q\,\mathrm{diag}(\sqrt{\sigma_{\mu,i}\sigma_{\nu,i}})\,Q^{\top}=J_{\infty}.$
Remark 4.2.
When the covariances commute (i.e., the Gaussian marginals share eigenvectors), OCD reduces to $n$ independent scalar Riccati flows, each converging to the correct cross–variance $\sqrt{\sigma_{\mu,i}\sigma_{\nu,i}}$. If the covariances do not commute, the symmetry of $J_{t}$ is preserved, preventing convergence to the generally non-symmetric OT solution.
Finally we address the speed of convergence to the OT solution in the case of commutting covariance matrices.
Theorem 4.3 (Exponential convergence of OCD to OT in the commuting Gaussian case).
Let $\mu=\mathcal{N}(0,\Sigma_{\mu})$ and $\nu=\mathcal{N}(0,\Sigma_{\nu})$ with $\Sigma_{\mu},\Sigma_{\nu}\succ 0$ and $\Sigma_{\mu},\Sigma_{\nu}]=0$. Consider the $L^{2}$-OCD dynamics initialized with independent samples, and let $J_{t}=\mathbb{E}[X_{t}Y_{t}^{\top}]$ denote the cross–correlation. Let
$J_{\textrm{OT}}\;=\;\Sigma_{\mu}^{1/2}\big(\Sigma_{\mu}^{1/2}\Sigma_{\nu}\Sigma_{\mu}^{1/2}\big)^{1/2}\Sigma_{\mu}^{1/2}$
be the Gaussian OT cross–correlation. Then $J_{t}\to J_{\textrm{OT}}$ exponentially fast: there exist constants $C>0$ and $\kappa>0$ such that
$\|J_{t}-J_{\textrm{OT}}\|_{2}\;\leq\;C\,e^{-2\kappa t}\qquad\text{for all }t\geq 0,$
where $\|\cdot\|_{2}$ denotes the spectral norm. Expressing $\Sigma_{\mu}=Q\operatorname{diag}(\sigma_{\mu,i})Q^{\top}$ and $\Sigma_{\nu}=Q\operatorname{diag}(\sigma_{\nu,i})Q^{\top}$ in a common orthonormal basis,
$\kappa\;=\;\min_{1\leq i\leq n}\Big(\sqrt{\tfrac{\sigma_{\nu,i}}{\sigma_{\mu,i}}}+\sqrt{\tfrac{\sigma_{\mu,i}}{\sigma_{\nu,i}}}\Big)\;\geq\;2.$
Proof.
Since $\Sigma_{\mu},\Sigma_{\nu}]=0$, there exists orthogonal $Q$ with $\Sigma_{\mu}=Q\operatorname{diag}(\sigma_{\mu,i})Q^{\top}$ and $\Sigma_{\nu}=Q\operatorname{diag}(\sigma_{\nu,i})Q^{\top}$. The Riccati ODE
$\dot{J}_{t}=\Sigma_{\mu}+\Sigma_{\nu}-J_{t}^{\top}\Sigma_{\mu}^{-1}J_{t}-J_{t}\,\Sigma_{\nu}^{-1}J_{t}^{\top}$
preserves symmetry, so writing $J_{t}=Q\operatorname{diag}(j_{i}(t))Q^{\top}$ gives $n$ decoupled scalar ODEs
$\dot{j}_{i}(t)=\sigma_{\mu,i}+\sigma_{\nu,i}-(\sigma_{\mu,i}^{-1}+\sigma_{\nu,i}^{-1})\,j_{i}(t)^{2}=\alpha_{i}\big((j_{i}^{\star})^{2}-j_{i}(t)^{2}\big),$
where $\alpha_{i}:=\sigma_{\mu,i}^{-1}+\sigma_{\nu,i}^{-1}$ and $j_{i}^{\star}:=\sqrt{\sigma_{\mu,i}\sigma_{\nu,i}}$. The explicit solution is
$j_{i}(t)=j_{i}^{\star}\,\tanh\!\big(\alpha_{i}j_{i}^{\star}t\big).$
Hence
$|j_{i}(t)-j_{i}^{\star}|=j_{i}^{\star}\Big|\,\tanh(\alpha_{i}j_{i}^{\star}t)-1\Big|=\frac{2j_{i}^{\star}\,e^{-2(\alpha_{i}j_{i}^{\star}t)}}{1+e^{-2(\alpha_{i}j_{i}^{\star}t)}}\;\leq\;2j_{i}^{\star}\,e^{-2\rho_{i}t},$
with $\rho_{i}:=\alpha_{i}j_{i}^{\star}=\sqrt{\sigma_{\nu,i}/\sigma_{\mu,i}}+\sqrt{\sigma_{\mu,i}/\sigma_{\nu,i}}$. Therefore, for the spectral norm we have
$\|J_{t}-J_{\textrm{OT}}\|_{2}=\big\|Q\,\operatorname{diag}(j_{i}(t)-j_{i}^{\star})\,Q^{\top}\big\|_{2}\leq\max_{i}|j_{i}(t)-j_{i}^{\star}|\leq 2\max_{i}j_{i}^{\star}\,e^{-2(\min_{i}\rho_{i})t}.$
Setting $C:=2\max_{i}j_{i}^{\star}$ and $\kappa:=\min_{i}\rho_{i}$ gives the claim. Finally, $\rho_{i}\geq 2$ (by arithmetic-geometric mean inequality), hence $\kappa\geq 2$. ∎
@@ 5. Numerical Algorithm through Opinion Dynamics
Opinion dynamics at its core tries to answer how different agents shape their opinions in contact with each-other [kolarijani2020macroscopic]. While seemingly unrelated to the Monge-Kantorovich problem, in this section, we leverage analogy between opinion dynamics and OCD to provide a novel, yet simple, numerical algorithm for the latter. The main idea is to first, split the data points into a certain number of clusters. Next, the conditional expectation is estimated in each cluster using an appropriate hypothesis class.
@@ 5.1. Analogy with Opinion Dynamics
Consider a population of agents indexed by $i\in\{1,...,N\}$. Suppose each person $i$ has certain two-dimensional position $P_{i}(t)\in\mathbb{R}^{2}$ as their opinion in time $t$. In the standard form of the bounded interval Hegselmann-Krause consensus dynamics, e.g., see [hegselmann2006truth, motsch2014heterophilious], these positions relax towards a global consensus via
@@
    The positions are perfectly tied to each-other and $\alpha$ becomes the exchange matrix $J$ (with ones on its anti-diagonal and zeros elsewhere).
@@
$\dot{P}_{X,i}=\frac{1}{N_{\mathrm{bar}{I}^{X_{i},\epsilon}}}\sum_{j\in\mathrm{bar}{I}^{X_{i},\epsilon}}(P_{Y,i}(t)-P_{Y,j}(t))\ \ \textrm{and}\ \ \dot{P}_{Y,i}=\frac{1}{N_{\mathrm{bar}{I}^{Y_{i},\epsilon}}}\sum_{j\in\mathrm{bar}{I}^{Y_{i},\epsilon}}(P_{X,i}(t)-P_{X,j}(t))\ ,$
@@
which is OCD (14) for $L^{2}$-cost. From technical point of view, the discrete version is justified due to the approximation of conditional expectation by a discrete kernel [nadaraya1964estimating]. Through the lens of opinion dynamics, the OCD describes a dynamics which is local in a sense that each agent is only influenced by agents in its vicinity, yet as discussed before, it is closely linked to the global optimization problem (2).
@@ 5.2. Non-parametric Monte-Carlo Algorithm
Starting with $N_{p}$, i.i.d., samples $\hat{X}_{i,t_{0}}\sim\mu$ and $\hat{Y}_{i,t_{0}}\sim\nu$, we update the state of each sample according to discretized form of OCD. It is important to note that here we are focusing on an approximate solution of the optimization problem (2) over atom-less distributions $\mu$ and $\nu$. An alternative interpretation, where the distributions are assumed to be sum of diracs (assignment problem) is not pursued here (see [trigila2016data]). Therefore throughout what follows, we assume that there exists an underlying continuous map between the two measure spaces. Three hyper-parameters control the discretization error: number of samples $N_{p}$, time step size $\Delta t$, and cluster cut-off $\epsilon$. Once these values are fixed, relying on Euler’s scheme (see Runge-Kutta’s scheme in Supplementary Information), each sample is updated according to
@@
    OCD-piecewise constant. The estimator is constant in each cluster (equivalent to using square kernel in the estimation of the conditional expectation). The expressions for $\hat{\mathcal{K}}^{n}_{X,Y}(.)$ are given in Supplementary Information.
@@
    OCD-piecewise linear. The estimator follows linear regression (accounts for using Gaussian kernel in the estimation of the conditional expectation). In other words here we look for $L^{2}$ projection of $\nabla_{x,y}c$ onto the span of linear functions [klebanov2021linear]. The final expressions are given in Supplementary Information.
The overview of the algorithm is given in Supplementary Information. The convergence of the algorithm, in general, is controlled by $N_{p},\Delta t,\epsilon$. Assuming that the OCD initialized by samples of $\mu,\nu$ is well-posed, it is reasonable to consider fixed-order convergence with respect to $\Delta t$ (e.g. first-order for Euler time integration, see [petzold1983automatic]). Also the error due to finite $N_{p}$ is expected to behave similar to conventional Monte-Carlo schemes [kroese2013handbook]. However, the error dependency on the cut-off value $\epsilon$ is more delicate. Two asymptotic limits of $\epsilon\to 0$ and $\epsilon\to\infty$ can be intuitively understood. The limit $\epsilon\to 0$ accounts for no interaction between agents, i.e. each point forms its own cluster. Therefore the points remain frozen in time. On the other hand, $\epsilon\to\infty$ gives rise to perfect interaction among all agents, forming one giant cluster. In this case, the dynamics converges to the average value among all sample points, for the 1st order approximation, and towards linear regression, for the 2nd order case. Obviously, the optimal value of $\epsilon$ needs to be tuned in-between, and is expected to depend on $N_{p}$. The transition between frozen dynamics in $\epsilon\to 0$ and homogenization $\epsilon\to\infty$ suggests a phase transition type behaviour, resulting from $\epsilon$. The transition point can serve as a proxy for the optimal cut-off value, as discussed in Supplementary Information.
[…]
@@ 5.3. Computational Complexity
In the following, we provide complexity estimates of OCD. Note that the discussion is mainly based on heuristic arguments and rely on practical scaling observations rather than formal asymptotic analysis. For a given time step (iteration), the main computational task of the proposed OCD algorithm is the evaluation of the conditional expectation, whose estimation relies on forming clusters by finding neighbours of each point in a given distribution. For the latter, though non-exclusively, the Ball Tree method [omohundro1989five] scales efficiently i.e. $\mathcal{O}(nN_{p}\log N_{p})$ with number of particles $N_{p}$ and space dimension $n$. Therefore, we expect that the OCD equipped with piecewise constant estimator for the conditional expectation to have the time complexity of $\mathcal{O}(nN_{p}\log N_{p})$, since the number of ODEs is $2n$ and number of operations in order to compute the distance between each two particles is $\mathcal{O}(n)$. However, since we deployed the direct solver in the computation of the conditional expectation with piecewise linear estimator, the corresponding OCD approach scales poorly with space dimension $n$ leading to the overall time complexity of $\mathcal{O}(n^{3}N_{p}\log N_{p})$.
In terms of memory consumption, since OCD, unlike linear programming, does not require computation of the distance matrix, its memory requirement is linear with the number of particles (histogram-dimension) $N_{p}$. In terms of space dimension $n$, OCD becomes memory intensive at high $n$ since the deployed direct linear solver scales poorly in the high dimensional setting.
In order to estimate the total computational cost of OCD, including number of iterations required to reach stationary state, we conducted numerical tests as shown in Fig. 2, where OCD runs between two normal distributions. The flops and storage requirements confirm our analysis on complexity and memory consumption. However, it is evident that OCD requires significantly more operations compared to the reference Earth Mover’s Distance (EMD) method [JMLR:v22:20-451], if relatively small number of sample points (e.g. 100) is considered. This is mainly due to the fact that while the number of operations scales efficiently, i.e. with $\mathcal{O}(N_{p}\log N_{p})$, the pre-factor which depends on the number of iterations required to reach stationary condition becomes dominant, if $N_{p}$ is relatively small. For $N_{p}>64^{\prime}000$, we observe that OCD becomes more efficient than EMD. On the other hand, as expected, OCD with the piecewise linear approximation of the conditional expectation scales poorly with the dimension, i.e. $\mathcal{O}(n^{3.5})$, while the measured timing of OCD with the piecewise constant estimator of the conditional expectation is slightly larger than expected $\mathcal{O}(n)$. We performed this study using the Google’s Cloud TPU v6e with 172.9 GB RAM.
@@
(a) $n=3$ (b) $N_{p}=1000$
Figure 2. Execution time (top) and memory consumption (down) of OCD-piecewise linear algorithm with Runge Kuta 4th order solver with $\Delta t=0.1$ against Earth Mover’s Distance (EMD) method [JMLR:v22:20-451] in learning the map between samples of $\mu=\mathcal{N}(0,I)$ and $\nu=\mathcal{N}(1,I)$ in (a) $n=3$ for a range of $N_{p}$ and (b) n=1,,128n=1,...,128 with $N_{p}=1000$. Here we scale $\epsilon$ with a rule of thumb $\varepsilon=3n{N_{p}}^{-1/4}/4$.